OpenStudy (anonymous):

[lim h-> 0] ((x+h)^2 -3 (x+h) - x^2+3x)/h how do i begin to find the limit? help please

OpenStudy (anonymous):

First off evaluate your expression

OpenStudy (anonymous):

Simplify (x+h)^2 -3 (x+h) - x^2+3x

OpenStudy (anonymous):

\[(x+h)^2 -3 (x+h) - x^2+3x\] \[=(x^2+2xh+h^2 -3x-3h- x^2+3x)/h\] \[=(2xh+h^2 -3h)/h\]

OpenStudy (anonymous):